OpenStudy (anonymous):

The slope of a line is 0, and the y-intercept is 6. What is the equation of the line written in slope-intercept form? is it y = 6x

OpenStudy (anonymous):

just 6

OpenStudy (agent47):

if the slope is 0, it's a horizontal line

OpenStudy (anonymous):

so y=6?

OpenStudy (anonymous):

yep

OpenStudy (agent47):

aka y-6=0(x-0) y-6=0 y=6